# Improving Diffusion Language Model Decoding through Joint Search in Generation Order and Token Space

### v1

@@ 1 Introduction
Diffusion Language Models (DLMs) have emerged as a promising alternative to autoregressive models for sequence generation. In particular, Masked Diffusion Models (MDMs) (Sahoo et al., 2024; Shi et al., 2024) trains on a core objective: learning to reconstruct original text by denoising sequences corrupted with random masks at varying rates. At inference, generation starts from a fully masked sequence and proceeds over multiple denoising steps: the model predicts token distributions for all positions, samples tokens for masked positions, and re-mask a subset of tokens to construct the next input. One key opportunity is that MDMs enable order-agnostic decoding: they revise many positions in parallel, allowing adaptive generation orders rather than left-to-right. This flexibility makes the choice of generation order a central design decision for decoding.
@@
Empirically, we find that existing methods occupy opposite ends of an exploration–exploitation trade-off. We diagnose exploration using pass@$k$, the probability that at least one of $k$ samples is correct. Low-confidence remasking exploits by committing early to high-confidence tokens and restricting decoding to a narrow set of trajectories, which boosts pass@$1$ but leads to slow pass@$k$ growth as $k$ increases. In contrast, random remasking explores a much broader set of trajectories and attains higher pass@$k$, but achieves weak pass@$1$ because it cannot reliably select good ones. Thus, prior methods either over-exploit or explore blindly, leaving much of the model’s multi-trajectory potential unrealized. We address this with explicit trajectory search and pruning.
@@ 2 Background
This section establishes the technical foundation for our work. We review the fundamentals of MDMs, formalize key concepts for remasking strategies, and define our evaluation metrics.
@@ 2.1 Discrete Diffusion Models
Discrete diffusion models adapt the forward diffusion process and the reverse denoising process (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song and Ermon, 2019; Song et al., 2021) to discrete data by establishing the diffusion process over a discrete domain $\mathbf{x}\in\mathcal{X}$, where $\mathbf{x}$ is a one-hot vector denoting tokens from a vocabulary of size $|\mathcal{X}|$ (Austin et al., 2021). Given a prior $\bm{\pi}$, the forward process $q$ incrementally corrupts the original data $\mathbf{x}_{0}$ into a target prior distribution $\text{Cat}(\cdot;\bm{\pi})$. Over continous time $t\in[0,1]$, it forms a sequence of increasingly noisy latent variables $\mathbf{x}_{t}$, through the conditional marginal distribution $q(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\text{Cat}\left(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+(1-\alpha_{t})\bm{\pi}\right)$. Here, $\alpha_{t}$ is a monotonically decreasing noise schedule that satisfies boundary conditions $\alpha_{0}=1$ and $\alpha_{1}=0$. Furthermore, we can achieve the transition probability between any two intermediate time points $0<s<t<1$ through $q(\mathbf{x}_{t}\mid\mathbf{x}_{s},\mathbf{x}_{0})=\text{Cat}\left(\mathbf{x}_{t};\alpha_{t}/\alpha_{s}\mathbf{x}_{s}+(1-\alpha_{t}/\alpha_{s})\bm{\pi}\right)$.
MDM, a specific instance of this framework, utilizes the prior $\bm{\pi}=\bm{m}$ to achieve absorbing-state diffusion, a particularly suitable setting for language modeling (Sahoo et al., 2024; Shi et al., 2024; Lou et al., 2024). Here, $\bm{m}$ is a one-hot vector corresponding to a special MASK token. Defining $s$ as the time step immediately preceding $t$, the posterior distribution simplifies to:
@@
2.2 Sampling Strategies in MDM
In masked generative models, sampling starts from a fully masked sequence, $\mathbf{x}_{1}=(\text{MASK},\ldots,\text{MASK})$. The model then iteratively refines this sequence over a series of steps. At each step, the model predicts logits for all currently masked tokens. The critical action in this reverse process is the transfer of a prediction—that is, the act of replacing a selected MASK token with its predicted value, thereby committing to that prediction for subsequent steps. The rule that determines which masked token to transfer next is known as the remasking strategy, and it defines the decoding order. We focus on three primary strategies:
Random Remasking. The strategy used during training. The next token to unmask is chosen uniformly at random from the set of all remaining masked tokens. This is a baseline that ensures unbiased, order-agnostic generation. Autoregressive (AR). We force the MDM to keep the leftmost predicted token and remask all following tokens. This baseline decouples the effect of generation order and solely examines the effect of diverse token selection. Low-Confidence Remasking. A common inference-time strategy. The token with the highest predicted probability is unmasked next; the tokens with lower probability are remasked. Formally, at each step, the model computes a confidence score for each masked token $i$ as its maximum logit, $s_{i}=\max(p_{\theta}(\cdot\mid\mathbf{x}_{t})_{i})$. The token with the maximum score $s_{i}$ is transferred. The intuition is to resolve the token position where the model has the greatest certainty first, potentially mitigating error propagation (Nie et al., 2025; Kim et al., 2025a).
2.3 Evaluation Metrics for Reasoning Performance
@@
The gap between pass@$1$ and pass@$k$ diagnoses how much a decoding strategy benefits from sampling. If pass@$1$ is strong but pass@$k$ increases slowly with $k$, the generated samples are highly correlated, indicating conservative decoding that commits early and explores few distinct trajectories. Conversely, if pass@$1$ is weak but pass@$k$ rises quickly with $k$, the model can reach correct solutions under sampling, but the decoding strategy cannot reliably select them. Our goal is to translate this sampling-driven capability into higher pass@$1$ by explicitly exploring and pruning in the joint space of generation orders and token choices.
Figure 2: Decoding strategies exhibit an exploration–exploitation trade-off. Empirical pass@$k$ curves for LLaDA-8B-Instruct and LLaDA-1.5 on reasoning and coding benchmarks. Low-confidence remasking often attains higher pass@$1$, but yields slow pass@$k$ growth as $k$ increases. In contrast, random remasking consistently, and autoregressive (AR) decoding in several settings, starts from lower pass@11 yet achieves substantially higher pass@$k$ at large $k$ ($\approx$256), reflecting broader coverage of the solution space through more diverse trajectories. These curves highlight an opportunity for decoding algorithms that translate high-pass@$k$ potential into stronger pass@11 by explicitly exploring and selecting among diverse trajectories.
3 Trade-off in Standard Decoding Strategies
Decoding Strategy Trade-off. To understand the trade-off in DLM decoding strategies, we investigate how different approaches affect reasoning performance by addressing the question: How does the diversity of generation trajectories explored by a decoding strategy relate to its ability to solve complex problems? We analyze three core strategies representing distinct exploration-exploitation trade-offs: random remasking (maximizing generation order diversity), low-confidence remasking (greedily exploiting local confidence), and fixed autoregressive order (enforcing generation order while enabling diverse token selection). This comparison is particularly valuable because DLMs offer unique flexibility in generation order compared to autoregressive models, yet optimal strategies for leveraging this flexibility remain unclear.
@@ 3 Trade-off in Standard Decoding Strategies
In this section, we study how exploration in decoding affects reasoning in MDMs. Specifically, we ask how the diversity of sampled decoding trajectories impacts problem-solving performance. We compare three representative strategies that emphasize different sources of diversity: low-confidence remasking (greedy, low diversity), random remasking (high diversity in generation order), and Autoregressive decoding (high diversity in token choices). This comparison isolates whether gains come from flexible order itself or from broader exploration in the joint space of order and token choices.
@@
Low-confidence remasking: high single-sample accuracy but rapid performance plateau. Figure 2 presents the key findings from our analysis. Using LLaDA models (Nie et al., 2025; Zhu et al., 2025) trained for flexible-order generation, we evaluate each strategy’s single-sample accuracy (pass@$1$) and multi-sample coverage (pass@$k$) across mathematical reasoning (GSM8K, MATH500) and coding (HumanEval, MBPP) benchmarks to study the effects of generation order and token selection diversity on reasoning performance. As expected, the low-confidence remasking strategy achieves the highest pass@$1$ in most cases, with advantages of 0.2-0.3 on MBPP. This aligns with the intuition that resolving the most certain tokens first helps guide the generation and reduces error propagation in a single sample. However, as we increase the sample budget $k$, the performance of low-confidence remasking plateaus relatively quickly.
Diverse strategies: lower initial accuracy but superior multi-sample coverage. In contrast, both random remasking and fixed AR order strategies, while starting at lower pass@$1$, continue to improve with more samples, eventually achieving significantly higher pass@$k$ (e.g., random/AR strategies reach $\sim$0.8 while low-confidence plateaus at $\sim$0.4 at pass@$256$ on MATH500). This demonstrates that diverse generation orders (from random remasking) and token selections (from AR decoding with temperature) can enable the model to solve more unique problems. We also conducted an experiment (Appendix A.2) showing that AR is not fundamentally superior than other generation orders. This implies that exploration in the generation order and token space, rather than the specific left-to-right order, is what improves the model’s reasoning performance. The sets of solutions generated through these diverse trajectories have greater coverage of the solution space, even though each individual sample is less reliable.
This consistent phenomenon reveals a fundamental limitation of greedy decoding strategies: by committing early to specific token choices based on local confidence, low-confidence remasking restricts exploration of the joint space of generation orders and token selections. While effective for optimizing a single sample, this approach fails to utilize the model’s full reasoning potential. The model’s inherent capability, as measured by pass@$k$ with large $k$, is better realized by strategies that introduce more stochasticity.
Our analysis reveals the core opportunity in DLM decoding: achieving higher accuracy than low-confidence remasking requires strategic exploration of both generation order and token space. To this end, we introduce Order-Token Search, an algorithm that actively searches this joint space within a single decoding process to locate correct answers that greedy strategies miss.
@@ 4 Method: Order-Token Search for Diffusion Language Models
The empirical results in Section 3 establish the need for a decoding algorithm that explores multiple potential generation trajectories. Standard one-trial sampling is insufficient as it commits to a single greedy sequence. To address this, we develop Order-Token Search, a novel algorithm inspired by beam search but tailored for the parallel, iterative nature of DLMs. The key innovation is a joint search over both the token selections and the order in which they are generated, guided by the model’s own likelihood predictions to score and prune candidate trajectories. This section details its two key components—search and prune. The complete algorithm is provided in Appendix A.1.
@@ 4.1 Search Process
We begin with $K$ identical copies (beams) of the initial sequence $\mathbf{x}_{1}=[\mathbf{c};\text{MASK}^{L}]$, where $\mathbf{c}$ is the prompt and $\text{MASK}^{L}$ denotes $L$ mask tokens. Over continuous denoising time $t\in[0,1]$, we independently apply the MDM to each candidate, generating new hypotheses that represent different choices in the joint space of tokens and generation orders. Between any two user-specified time $(s,t)$, our algorithm can perform search and expand each candidate to become multiple candidates independently with additional denoise steps. This expansion explores the joint space of tokens and generation orders, creating a diverse set of candidate sequences which are then evaluated and pruned to retain the top-$K$ trajectories with the highest model likelihood (see Section 4.2).
To manage computational complexity, we structure the search using block diffusion (Arriola et al., 2025; Nie et al., 2025). Instead of searching at every denoising step—which would incur an $O(K\cdot|t|)$ overhead for $|t|$ steps—we perform the search expansion only at the boundaries between contiguous blocks of tokens. This reduces the overhead to $O(K\cdot|b|)$, where $|b|$ is the total number of blocks, making the search tractable. After processing all blocks, the single best sequence is selected from the final $K$ candidates based on the highest likelihood. Details on computational complexity are provided in Appendix A.4.
@@ 4.2 Pruning Process
The effectiveness of Order-Token Search hinges on a pruning criterion that can accurately score candidate sequences with diverse tokens and generation orders. Our key insight is that the standard MDM objective can be an unreliable scorer, as the model is trained on—and often fails at—extremely difficult infilling tasks where a large number of tokens must be predicted simultaneously (Kim et al., 2025a). To obtain a more stable and accurate likelihood estimate, we instead score a candidate based on the incremental denoising actions that created it.
We propose a scoring function $s$ that evaluates the model’s confidence for each discrete denoising step. For a step from a more corrupted state $\bm{x}_{t}$ to a less corrupted state $\bm{x}_{s}$ (where $0\leq s<t\leq 1$), the score is calculated as:
$s(\bm{x}_{t};\bm{x}_{s})=\mathbb{E}_{\bm{x}_{0}\sim p_{\theta}(\bm{x}_{0}\mid\bm{x}_{t})}\log p(\bm{x}_{0}|b(\bm{x}_{s},\bm{x}_{t},\bm{x}_{0})),$ (2)
where $p(\cdot|\cdot)$ is the parametrized posterior from Section 2.1. The function $b$ identifies the specific blocks of token positions $\{i\mid\bm{x}_{t,i}=\text{MASK}\cap\bm{x}_{s,i}\neq\text{MASK}\}$ that were denoised between time $t$ and $s$, masks these blocks in $\bm{x}_{0}$, and returns the masked sequence. The score $s(\bm{x}_{t};\bm{x}_{s})$ is the log-likelihood of only these newly-revealed blocks, conditioned on the surrounding context provided by the model’s full-sequence prediction $\bm{x}_{0}$.
@@
This approach provides a better likelihood estimation with lower variance. By focusing on smaller, incremental predictions, we assess the model on tasks similar to its well-learned training distribution, where it denoises a limited number of masks at a time. The total score for a candidate sequence is the sum of scores over all its search-guided denoising steps: $\sum_{(s,t)\in\mathcal{I}}s(\bm{x}_{t};\bm{x}_{s})$, where $\mathcal{I}$ is the set of intervals where a search was performed. This sum captures the entire history of the candidate’s generation trajectory. Figure 3 illustrates this process for a single step.
In summary, Order-Token Search performs a joint search over tokens and generation orders, guided by the MDM’s own likelihood. By allowing candidates to explore denoising trajectories independently, Order-Token Search achieves diverse and effective exploration of the joint output space. The algorithm evaluates this exploration by scoring a candidate’s entire generation history through incremental block-level likelihoods, providing a comprehensive measure of global coherence. This approach efficiently leverages the iterative, parallel nature of MDMs: block diffusion minimizes computational overhead, while the full-sequence prediction $\bm{x}_{0}$ supplies rich context for stable likelihood estimation. Consequently, Order-Token Search enables effective pruning of low-likelihood trajectories, steering the search toward high-quality, coherent outputs.

### v2

@@ 1 Introduction
Diffusion Language Models (DLMs) have emerged as a promising alternative to autoregressive models for sequence generation. In particular, Masked Diffusion Models (MDMs) (Sahoo et al., 2024; Shi et al., 2024) trains on a core objective: learning to reconstruct original text by denoising sequences corrupted with random masks at varying rates. At inference, generation starts from a fully masked sequence and proceeds over multiple denoising steps: the model predicts token distributions for all positions, samples tokens for masked positions, and re-masks a subset of tokens to construct the next input. One key opportunity is that MDMs enable order-agnostic decoding: they revise many positions in parallel, allowing adaptive generation orders rather than left-to-right. This flexibility makes the choice of generation order a central design decision for decoding.
@@
Empirically, existing methods occupy opposite ends of an exploration–exploitation trade-off. The pass@$k$ metric captures the probability that at least one of $k$ samples is correct. Low-confidence remasking exploits by committing early to high-confidence tokens and restricting decoding to a narrow set of trajectories, which boosts pass@$1$ but leads to slow pass@$k$ growth as $k$ increases. In contrast, random remasking explores a much broader set of trajectories and attains higher pass@$k$, but achieves weak pass@$1$ because it cannot reliably select good ones. Thus, prior methods either over-exploit or explore blindly, leaving much of the model’s multi-trajectory potential unrealized. We address this with explicit trajectory search and pruning.
@@ 2 Background
This section establishes the technical foundation for our work. We review fundamentals of Masked Diffusion Models, formalize key concepts for remasking strategies, and define our evaluation metrics for the exploration trade-off.
@@ 2.1 Discrete Diffusion Models
Diffusion Language Models extend discrete diffusion modeling to text, and recent large-scale training has made them competitive with autoregressive models (Nie et al., 2025; Ye et al., 2025). Concretely, discrete diffusion models adapt the forward diffusion process and the reverse denoising process (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song and Ermon, 2019; Song et al., 2021) to discrete data by establishing the diffusion process over a discrete domain $\mathbf{x}\in\mathcal{X}$, where $\mathbf{x}$ is a one-hot vector denoting tokens from a vocabulary of size $|\mathcal{X}|$ (Austin et al., 2021). Given a prior $\bm{\pi}$, the forward process $q$ incrementally corrupts the original data $\mathbf{x}_{0}$ into a target prior distribution $\text{Cat}(\cdot;\bm{\pi})$. Over continous time $t\in[0,1]$, it forms a sequence of increasingly noisy latent variables $\mathbf{x}_{t}$, through the conditional marginal distribution $q(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\text{Cat}\left(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+(1-\alpha_{t})\bm{\pi}\right)$. Here, $\alpha_{t}$ is a monotonically decreasing noise schedule that satisfies boundary conditions $\alpha_{0}=1$ and $\alpha_{1}=0$. Furthermore, we can achieve the transition probability between any two intermediate time points $0<s<t<1$ through $q(\mathbf{x}_{t}\mid\mathbf{x}_{s},\mathbf{x}_{0})=\text{Cat}\left(\mathbf{x}_{t};\alpha_{t}/\alpha_{s}\mathbf{x}_{s}+(1-\alpha_{t}/\alpha_{s})\bm{\pi}\right)$.
Masked Diffusion Model (MDM), a specific instance of this framework, utilizes the prior $\bm{\pi}=\bm{m}$ to achieve absorbing-state diffusion, a particularly suitable setting for language modeling (Sahoo et al., 2024; Shi et al., 2024; Lou et al., 2024). Here, $\bm{m}$ is a one-hot vector corresponding to a special MASK token. Defining $s$ as the time step immediately preceding $t$, the posterior distribution simplifies to:
@@
2.2 Remasking Strategies in MDM
In masked generative models, decoding starts from a fully masked sequence, $\mathbf{x}_{1}=(\text{MASK},\ldots,\text{MASK})$. The model then iteratively refines this sequence over a series of steps. At each step, the model predicts logits for all currently masked positions. The critical action in this reverse process is the transfer of a prediction—that is, the act of replacing a selected MASK token with its predicted value, thereby committing to that prediction for subsequent steps. The rule that determines which masked token to transfer next is known as the remasking strategy, and it defines the decoding order. We focus on three primary strategies:
Random Remasking. The strategy used during training. The next position to unmask is chosen uniformly at random from the set of all remaining masked positions. This is a baseline that ensures unbiased, order-agnostic generation. Low-Confidence Remasking. A common inference-time strategy. The position with the highest confidence score is unmasked next; the positions with lower scores are remasked. Formally, at each step, the model computes a confidence score for position $i$ as the predicted probability assigned to the currently selected token at that position $s_{i}=p_{\theta}(\cdot\mid\mathbf{x}_{t})_{i}$. The position with the maximum score $s_{i}$ is unmasked. The intuition is to resolve the token position where the model has the greatest certainty first, potentially mitigating error propagation (Nie et al., 2025; Kim et al., 2025a). Autoregressive (AR). We force the MDM to keep the leftmost predicted token and remask all following tokens. This baseline decouples the effect of generation order and solely examines the effect of diverse token selection.
2.3 Evaluation Metrics for Exploration Trade-off
@@
The gap between pass@$1$ and pass@$k$ measures how much a decoding method benefits from sampling. If pass@$1$ is strong but pass@$k$ increases slowly with $k$, the generated samples are highly correlated, indicating conservative decoding that commits early and explores few distinct trajectories. Conversely, if pass@$1$ is weak but pass@$k$ rises quickly with $k$, the model can reach correct solutions under sampling, but the decoding method cannot reliably select them. Our goal is to translate this sampling-driven capability into higher pass@$1$ by explicitly exploring and pruning in the joint space of generation orders and token choices.
Figure 2: Standard DLM decoders reveal an exploration–exploitation trade-off. Empirical pass@$k$ curves for LLaDA-8B-Instruct and LLaDA-1.5 on mathematical reasoning and coding benchmarks. Low-confidence remasking often achieves higher pass@$1$ by exploiting local confidence, but its pass@$k$ grows slowly with $k$, indicating limited coverage from a narrow set of trajectories. In contrast, random remasking (order diversity) and AR decoding (token diversity) typically start with lower pass@$1$ yet reach much higher pass@$k$ by $k=256$, reflecting broader solution-space coverage through more diverse trajectories. We use pass@$k$ here only as a diagnostic of latent multi-trajectory potential; the goal is to convert this potential into improved pass@11 via structured exploration and selection.
3 Trade-off in Standard Decoding Methods
@@
In this section, we study the exploration–exploitation trade-off among common DLM decoding methods. Specifically, we ask how the diversity of sampled decoding trajectories impacts problem-solving performance. We compare three representative approaches that emphasize different sources of diversity: low-confidence remasking (greedy, low diversity), random remasking (high diversity in generation order), and autoregressive decoding (high diversity in token selections). This comparison is particularly valuable because DLMs offer unique flexibility in generation order compared to autoregressive models, yet optimal strategies for leveraging this flexibility remain unclear.
@@
Low-confidence remasking: higher pass@$1$ but slower pass@$k$ growth. Figure 2 summarizes our key findings. Using LLaDA models (Nie et al., 2025; Zhu et al., 2025) trained for flexible-order generation, we evaluate each decoding method in terms of single-sample accuracy (pass@$1$) and multi-sample coverage (pass@$k$) on mathematical reasoning (GSM8K, MATH500) and coding (HumanEval, MBPP) benchmarks, to understand how diversity in generation order and token selection affects performance. Across 8 dataset–model settings, low-confidence remasking achieves the best pass@$1$ in 4 and the second-best in 3 (e.g., improving MBPP by 0.2–0.3 absolute). This competitive pass@$1$ performance is consistent with the intuition that committing first on high-confidence tokens provides stable context and reduces error propagation within a single trajectory. However, as the sampling budget $k$ increases, pass@$k$ for low-confidence remasking grows more slowly and is eventually overtaken by other baselines. This is likely because low-confidence remasking over-exploits local confidence and narrows the set of feasible decoding trajectories.
Exploratory approaches: generally lower pass@$1$ but faster pass@$k$ growth. In contrast, random remasking and left-to-right autoregressive (AR) order typically start from a lower pass@$1$, but benefit much more from additional samples: pass@$k$ keeps rising with $k$ and eventually surpasses low-confidence remasking across most settings (e.g., on MATH500 at pass@$256$, random/AR reach $\sim$0.8 while low-confidence ends near $\sim$0.4). This behavior suggests that diversity enables the model to uncover correct solutions that are missed by a single committed trajectory. Such diversity can arise either from varying generation orders (random remasking) or from varying token selections under a fixed order (AR decoding with temperature). We also observe that AR decoding performs best on mathematical reasoning, achieving the highest pass@$1$ and pass@$k$ on GSM8K and MATH500. A plausible explanation is that these datasets more closely follow left-to-right, step-by-step solution structure, making a fixed causal order advantageous. In contrast, this pattern does not extend to coding: on HumanEval, AR attains the lowest pass@$1$. Overall, these results underscore the importance of diversity in both generation order and token choices for achieving broad coverage across tasks.
Summary and implication. Standard decoding methods expose a clear exploration–exploitation trade-off in DLMs. Greedy methods such as low-confidence remasking exploit local confidence to optimize pass@$1$, but do so by collapsing decoding to a narrow set of trajectories, leaving many valid solutions unexplored. More exploratory approaches, in contrast, reveal the model’s ability to solve additional problems by exploring alternative generation orders or token choices, though this comes at the cost of lower per-sample reliability. Importantly, our use of pass@$k$ serves only to diagnose this latent potential for exploration—not as a target metric. The central challenge, therefore, is how to leverage the benefits of diversity to increase single-sample accuracy.
@@ 4 Method: Order-Token Search for Diffusion Language Models
Section 3 shows that improving pass@$1$ requires converting trajectory diversity into a reliable single sample. To this end, we introduce Order-Token Search, a novel algorithm inspired by beam search but designed for the parallel, order-agnostic dynamics of DLMs. Order-Token Search performs search in the joint space of generation order and token choices, and uses the model’s likelihood predictions to score partial trajectories and stably prune unpromising ones. This section presents its two key components—search and prune—with full pseudocode in Appendix A.1.
@@ 4.1 Search Process
Order-Token Search maintains a beam set of $K$ partially denoised candidates, initialized as $K$ identical copies of $\mathbf{x}_{1}=[\mathbf{c};\text{MASK}^{L}]$, where $\mathbf{c}$ is the prompt and $\text{MASK}^{L}$ denotes $L$ mask tokens. Decoding proceeds over continuous denoising time $t\in[0,1]$. Unlike standard one-trajectory sampling, which fixes a single sequence of remasking decisions, our goal is to explicitly branch over both sources of uncertainty: what tokens to write and when/where to write them (i.e., the induced generation order).
Concretely, at a user-specified interval $(s,t)$, we expand each beam independently for a number of denoising steps. During expansion, the MDM produces token distributions for all currently masked positions, and the algorithm proposes multiple denoising actions that correspond to different joint samples from the order-token space (e.g., different subsets of positions to update and different sampled token assignments at those positions). Each proposed action yields a new candidate state, thereby creating multiple partial trajectories from the same prefix. This branching step is key for Order-Token Search to explore diverse trajectories, while downstream pruning (Section 4.2) keeps only the top-$K$ candidates according to a stable likelihood estimate.
To manage computational complexity, we organize expansions using block diffusion (Arriola et al., 2025; Nie et al., 2025). Branching at every denoising step introduces an $O(K\cdot|T|)$ overhead over $|T|$ total steps. Instead, we only perform expansion at the boundaries of contiguous token blocks, and run independent denoising within each block. This reduces the number of decision points from $|T|$ to $|B|$, where $|B|$ is the total number of blocks, yielding $O(K\cdot|B|)$ search overhead while still allowing substantial diversity in both generation order and token values. After processing all blocks, we select a single output sequence from the remaining $K$ candidates based on the highest likelihood. A detailed complexity analysis is provided in Appendix A.3.
@@ 4.2 Pruning Process
The effectiveness of Order-Token Search depends on a pruning criterion that can reliably compare partially denoised candidates whose trajectories may differ in both token values and generation order. A direct choice is to score candidates using the standard MDM objective, i.e., the likelihood of reconstructing the full clean sequence from the current noisy state. However, this signal can be unstable in practice: when many tokens remain masked, the model is implicitly evaluated on extremely difficult infilling problems that are far from the “small-step” denoising behavior it learns well from training (Kim et al., 2025a), and these scores can become high-variance and poorly calibrated for ranking.
Scoring incremental denoising actions. Our key idea is to score a candidate by the specific denoising actions that produced it, rather than by a monolithic full-sequence likelihood at an intermediate state. Consider a transition from a more corrupted state $\bm{x}_{t}$ to a less corrupted state $\bm{x}_{s}$ with $0\leq s<t\leq 1$. Let $\bm{x}_{0}\sim p_{\theta}(\bm{x}_{0}\mid\bm{x}_{t})$ denote a sample from the model’s conditional distribution over clean sequences given $\bm{x}_{t}$. We define an action-level score
$s(\bm{x}_{t};\bm{x}_{s})=\mathbb{E}_{\bm{x}_{0}\sim p_{\theta}(\bm{x}_{0}\mid\bm{x}_{t})}\log p_{\theta}\left(\bm{x}_{0}\,\middle|\,b(\bm{x}_{s},\bm{x}_{t},\bm{x}_{0})\right),$ (2)
where $p_{\theta}(\cdot|\cdot)$ is the parametrized posterior from Section 2.1. The function $b$ extracts the newly revealed blocks between $t$ and $s$, i.e., $\{i\mid\bm{x}_{t,i}=\text{MASK}\ \cap\ \bm{x}_{s,i}\neq\text{MASK}\}$, then masks exactly these positions in $\bm{x}_{0}$ and returns the resulting masked sequence. Intuitively, $s(\bm{x}_{t};\bm{x}_{s})$ measures how confidently the model can re-predict only the tokens that were newly denoised in this step, using the surrounding context supplied by the model’s own full prediction $\bm{x}_{0}$. Figure 3 visualizes this block-level scoring for one transition.
@@
Trajectory-level score for pruning. This action-level view yields a lower-variance and more comparable signal across candidates. Because each term focuses on a small set of tokens, it avoids penalizing an intermediate state for being “globally hard” when many masks remain, and instead evaluates whether a trajectory’s past decisions (which order and what tokens it chose) are supported by the model. We assign each candidate a cumulative score by summing over the set of intervals $\mathcal{I}$ at which the algorithm performed branching and search: $S(\text{candidate})=\sum_{(s,t)\in\mathcal{I}}s(\bm{x}_{t};\bm{x}_{s}).$ This sum captures the full generation history of partial trajectories: candidates that repeatedly make model-consistent denoising moves accumulate higher scores, while trajectories that rely on unlikely commitments are pruned early.
Overall, Order-Token Search alternates between searching and pruning to turn trajectory diversity into improved accuracy (pass@$1$). The search process maintains partial trajectories sampled from a joint distribution over which positions to reveal (generation order) and what tokens to write (token values). The pruning process then ranks these candidates using a trajectory-level score that decomposes likelihood into incremental denoising moves while providing a comprehensive measure of global coherence. By repeatedly keeping only the top-$K$ candidates after each search expansion, Order-Token Search allocates a fixed inference-time compute budget toward exploring diverse order-token trajectories and reliably selecting the most coherent one.